Acceleron Inc. is planning to expand its present facilities. The company will make the move when its real estate sinking fund has a total value of P1.2 million. If the fund currently has P400,000 and the company adds P50,000 per year, how many years will it take for the account to reach the desired value, if it earns interest at a rate of 10% per year?
1
Expert's answer
2020-10-05T11:00:53-0400

The number of years to reach the desired value is:

"n = \frac{ln(1 + (1,200,000 - 400,000)×0.1/50,000) } {ln(1 + 0.1)} = \frac{ln(2.6)} {ln(1.1)} = 10.025"

or approximately 10 years.